Circular No. 8541

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2005ci IN NGC 5682 Further to IAUC 8534, D. R. Madison and W. Li report the LOSS

discovery of an apparent supernova on unfiltered KAIT images taken on June 10.31 (at mag 18.5) and 11.26 UT (mag 18.4). SN 2005ci is located at α = 14^{h}34^{m}44^{s}.88, δ = +48^{o}40′19′′.8 (equinox 2000.0), which is 0′′.6 west and 6′′.1 north of the nucleus of NGC 5682. KAIT images taken on May 27.30 (limiting mag ∼ 19.5) and June 2.31 (limiting mag ∼ 19.0) showed nothing at this position.

SUPERNOVA 2005cj IN ESO 114-14 Further to IAUC 8517, L. A. G. Monard reports his discovery of an

apparent supernova (mag ∼ 17.9 ± 0.3) on images taken on 2005 June 12.120 and 13.134 UT with a 0.30-m reflector. SN 2005cj is located at α = 1^{h}54^{m}43^{s}.94, δ = −62^{o}06′32′′.0 (equinox 2000.0), which is 26′′ east and 14′′ south of the nucleus of the galaxy ESO 114-14. Nothing is visible at this location on an image taken by Monard on 2005 May 11.14 (limiting red mag 18.5) or on an image from the Digitized Sky Survey (limiting red mag 20.5).

SUPERNOVA 2005ch M. Modjaz, R. Kirshner, and P. Challis, Harvard-Smithsonian Center

for Astrophysics; and T. Matheson, National Optical Astronomy Observatory, report that a spectrum (range 340–730 nm) of SN 2005ch (cf. IAUC 8539, 8540), obtained on June 10.28 UT by M. Calkins with the F. L. Whipple Observatory 1.5-m telescope (+ FAST), shows it to be a type-Ia supernova well before maximum light. The spectrum consists of a blue continuum and the characteristic Si II feature (rest 635.5 nm), in addition to absorption lines due to Ca II, Mg II, Fe II, Si II, and S II.

M. Dennefeld and F. Ricquebourg, Institut d’Astrophysique de Paris, write that a calibrated spectrum (range 465–830 nm, resolution 0.6 nm) of SN 2005ch, obtained on June 10.93 UT with the 1.93-m telescope (+ Carelec spectrograph) at Haute-Provence Observatory, shows a blue continuum, strong Si II (rest 635.5 nm) absorption feature (measured at 603.3 nm), and a well-developed Fe II 501-nm blend and S II doublet (547–565 nm) absorptions, which indicate a type-Ia supernova a few days before maximum. No feature is detected from the parent galaxy.

2005 June 12  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green